# BiTrajDiff: Bidirectional Trajectory Generation with Diffusion Models for Offline Reinforcement Learning

### v3

@@ 1 Introduction
Offline Reinforcement Learning (Offline RL) (fujimoto2019off) seeks to derive effective decision policies exclusively from pre-collected datasets, obviating the necessity for online exploration due to the inherent risks and costs associated with real-time interaction (levine2020offline). This paradigm has demonstrated notable efficacy across various real-world applications, including robotic control (qing2024a2po, qing2022XRLsurvey, liu2023curricular) and power grid management (chen2024powerformer, xu2024TPA). A primary challenge in offline RL is addressing the out-of-distribution (OOD) problem, stemming from the inherent divergence between the static offline dataset distribution and the learned policy behavioral distribution during training. Consequently, the direct application of conventional online RL methods inevitably leads to extrapolation error (levine2020offline), characterized by the overestimation of value functions for OOD state-action pairs. To tackle the OOD problem, offline RL research emphasizes the development of techniques that constrain learned policies to remain within the distribution of the static offline dataset. These approaches include imposing explicit constraint terms during policy optimization (qing2024a2po, chen2022lapo) and penalizing the value function estimates for OOD actions (lyu2022mildly, xu2023offline).
@@
Despite promising advancements in offline RL, current algorithms critically restrict agent actions, as evidenced by (levine2020offline, fujimoto2019off, qing2024a2po). This limitation stems from the inherent mechanism constraining learned policy to the biased distribution of offline datasets, often leading to overfitting on limited behavioral patterns. While this principle effectively facilitates the replication of favorable behavioral patterns within the data, it inevitably curtails generalizability, hindering the exploration of optimal behaviors beyond dataset coverage. A common approach to mitigate the distribution bias is Data Augmentation (DA) method, which enriches the offline dataset with a large amount of synthesized high-fidelity data. For instance, studies like (lu2023synthetic, li2024diffstitch, li2024augmenting) leverage one generative diffusion model to approximate the distribution of single-direction forward-future trajectories for augmenting offline data. Notably, as data augmentation (DA) methods operate by modifying the training dataset, they are orthogonal to specific offline RL algorithms, enabling them to enhance any such algorithm without altering its core optimization mechanisms. However, a critical limitation of DA methods is their exclusive focus on trajectories originating from a predefined initial state. While these generated single-direction trajectories effectively explore behaviors emerging from a given state, they neglect the exploration of trajectories leading to that state. This oversight restricts the utility of the generated trajectory to locally explore the trajectory space immediately surrounding existing data, specifically for transitions originating from states already in the dataset. Consequently, while DA methods successfully generate trajectories for local state pairs already reachable within the offline dataset, they are ineffective in discovering global trajectories for previously unreachable state pairs. This naturally leads us to a neglected yet unresolved question:
@@ 2 Related Works
Offline Reinforcement Learning encompasses four primary categories: policy constraint (wang2022diffusion, xu2023offline), value regularization (kostrikov2021offline, kumar2020conservative), model-based (yu2020mopo, chemingui2024offline), and return-conditioned supervised learning (chen2021decision, paster2022you). Policy constraint methods restrict policies to the offline dataset, employing techniques such as explicit behavior cloning regularization (qing2024a2po, chen2022lapo) or implicit optimization towards in-sample optimal policies (nair2020awac, yue2022boosting). Meanwhile, value regularization promotes conservative value functions to alleviate OOD overestimation via establishing precise lower bounds for Q-values to facilitate in-sample action selection (lyu2022mildly, park2024value). Model-based approaches construct task-oriented world models from offline data, including one-step dynamics (yu2020mopo, chemingui2024offline) and multi-step sequence models (hafner2020mastering, hafner2023mastering), which are subsequently used for offline reward penalization (yu2021combo, kidambi2020morel) and online decision planning to ensure in-distribution actions (hansen2022temporal, hansen2023td). Return-conditioned supervised learning trains trajectory generators conditioned on returns, with transformers (chen2021decision, schmied2024retrieval) and diffusion models (janner2022planning, ajay2022conditional) as common backbones. Despite their effectiveness, these approaches are fundamentally limited by reliance on the dataset distribution.
Data Augmentation for Offline RL synthesizes additional interactions to enrich the dataset for subsequent RL training. Early work employs dynamics ensembles to generate transitions with uncertainty weighting (zhang2023uncertainty) or diffusion models for flexible one-step oversampling (lu2023synthetic). More recent efforts shift toward trajectory-level generation. Some work focuses on on-policy rollouts, guided either by policy outputs (jackson2024policy) or interaction histories (he2023diffusion). Other approaches explore prioritized augmentation based on dynamics error or curiosity (wang2024prioritized), as well as trajectory recomposition through reverse generation (yang2025rtdiff) or diffusion-based stitching (li2024diffstitch). Despite these advances, existing DA methods remain limited by their forward-only generation paradigm, restricting exploration to local variations around the dataset. Our BiTrajDiff overcomes this limitation by jointly generating forward and backward trajectories from intermediate states.
@@ 3 Preliminary
Reinforcement Learning paradigm (sutton2018reinforcement) is formally defined as a Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma,\rho_{0}\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the environment dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\rho_{0}$ denotes the initial state distribution. At each timestep $t$, an agent executes an action $a_{t}$ according to its policy $\pi(\cdot|s_{t})$ in the current state $s_{t}$, transitioning to the subsequent state $s_{t+1}$ with reward $r_{t}$ based on $P$. The objective of the agent is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return: $\pi^{*}=\arg\max_{\pi}\mathcal{J}(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. In contrast to online RL, offline RL (levine2020offline) relys solely on a static trajectories dataset $\mathcal{D}=\{\tau^{i}\}_{i=1}^{N}$, where $N$ is the number of trajectories and each trajectory $\tau^{i}=\{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\}_{t=0}^{T-1}$ comprises a sequence of state-action-reward tuples.
Diffusion Models (ho2020denoising, song2021maximum) are generative frameworks that learn data distributions through a forward diffusion process and a reverse denoising process. The forward process gradually corrupts data $\mathbf{x}_{0}$ via a predefined noise schedule $\{\beta_{k}\}_{k=1}^{K}$, where each transition is defined as $q(\mathbf{x}_{k}|\mathbf{x}_{k-1})=\mathcal{N}\left(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}\right)$. As $K\to\infty$, the terminal distribution $q(\mathbf{x}_{K})$ converges to an isotropic Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process, parameterized by $\theta$, aims to iteratively reconstruct the original data through learnable Gaussian transitions: $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})=\mathcal{N}\left(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\Sigma_{\theta}(\mathbf{x}_{k},k)\right).$ By maximizing the Empirical Lower Bound (ELBO) (sohn2015learning) on the log-likelihood of the sampled data, the diffusion model can be trained with a simplified surrogate loss (sohn2015learning):
@@ 4.1 Bidirectional Diffusion Training
Building upon recent advances in diffusion-based sequence generation (fathi2025unifying, gong2022diffuseq), we propose a bidirectional trajectory generation framework comprising independent forward and backward trajectory diffusion models. These models are separately trained on offline datasets to learn the trajectory distribution of the behavior policy, facilitating bidirectional state trajectory generation from shared intermediate “anchor” states.
@@
where $\mathbf{x}_{0}(\tau)$ denotes the final generated subtrajectory of diffusion, $\mathbf{y}(\tau)$ represents the condition for generation, and $\theta$ represents the learnable parameters of the diffusion model. While various condition metrics are available, such as temporal difference error, Q-function (wang2024prioritized), our framework utilizes the cumulative return $R$ and the state $s$ as conditional input $\mathbf{y}(\tau)$. This approach, consistent with prior works (ajay2022conditional, ding2024diffusion), enables the generation of trajectories with desired returns starting from the specific states. However, to enable bidirectional trajectory generation, the formations of the backward and forward diffusion condition inputs are quite distinct.
@@
To realize conditional diffusion model training and inference, we employ the Classifier-Free Guidance (CFG) technique (ho2022classifier), which uses a single denoise model to handle both conditional and unconditional generation. Formally, the perturbed noise after CFG can be formed as:
@@
where $\text{Bern}(p)$ represents the Bernoulli distribution. In our implementation, we fix the state condition $s_{t}$ in $\mathbf{x}_{k}(\tau)$ during the adding-noise training as well as the denoising inference stage for simplicity and efficiency following (ajay2022conditional, yang2025rtdiff), while the diffusion denoise models $\epsilon_{\theta}$ only take the cumulative return $R(\tau)$ components of $\mathbf{y}(\tau)$ as condition inputs.
@@ 4.2.1 Bidirectional State Trajectory Generation
We utilize the pretrained forward and backward trajectory diffusion model in section 4.1 to generate bidirectional state trajectories. Prior methods (lu2023synthetic, li2024augmenting) model only the conditional distribution of trajectories originating from a given state, neglecting the historical pathways leading to those states. As a result, these data augmentation techniques yield only limited expansion in behavior diversity. To address this limitation without introducing additional noise, our proposed method, BiTrajDiff, generates both forward and backward trajectories and integrates them by stitching at a shared intermediate state $s_{t}$, which serves as a consistent anchor.
@@ 4.2.3 Bidirectional Trajectory Filter
In summary, our BiTrajDiff operates in two key stages. Initially, it trains a bidirectional diffusion model to capture the distributions of both forward-future and backward-history trajectories. Subsequently, BiTrajDiff iteratively generates bidirectional trajectory conditioned on intermediate anchor states sampled from $\mathcal{D}$. These generated sequences undergo further completion and filtering to create a high-fidelity augmented dataset $\tilde{\mathcal{D}}$, which is then mixed with $\mathcal{D}$ to train downstream offline RL algorithms. We adopt CleanDiffuser (dong2024cleandiffuser) as the backbone for BiTrajDiff, owing to its generality and robustness. The pseudocode of BiTrajDiff is presented in Appendix A.
@@ 5 Experiments
To demonstrate the effectiveness of the proposed BiTrajDiff method, we conduct experiments on the D4RL benchmark (fu2020d4rl). Our evaluation seeks to answer the following questions: (1) Does BiTrajDiff outperform existing DA methods across various offline RL algorithms? (Section 5.2 and Appendix C) (2) How do different components of BiTrajDiff affect the offline RL performance? (Section 5.3 and Appendix D-F) (3) Can BiTrajDiff find the potentially valuable state leading to high return more effectively than other DA baselines? (Section 5.4 and Appendix G) (4) Can BiTrajDiff generate more diverse trajectories compared to single-direction DA approaches? (Section 5.5)
@@ 5.1 Experiment Settings
We evaluate our approach on the D4RL benchmark (fu2020d4rl), focusing on locomotion, navigation, and manipulation tasks. We compare BiTrajDiff with three state-of-the-art DA methods: Synther (lu2023synthetic),DiffStitch (li2024diffstitch) and RTDiff (yang2025rtdiff). To evaluate the effectiveness of them, we conduct extensive experiments on four representative offline RL algorithms: IQL (kostrikov2021offline), TD3BC (fujimoto2021minimalist), CQL (kumar2020conservative), and DT (chen2021decision). These methods have been widely adopted as standard baselines in offline RL due to their stable performance. The hyperparameter setting can be viewed in Appendix B.2.
@@
Source Task  IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
@@
Task  IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
@@ 5.3 Ablation Analysis
Ablation Study of the Direction of Diffusion Generation. To confirm the effectiveness of our bidirectional generation paradigm, we compare the performance improvement with the augmented diffusion trajectories generated in a single direction and in a bidirectional way, as shown in Figure 2. While the forward method yields moderate improvements, it is consistently outperformed because it neglects the historical transitions leading to critical, high-reward states. Conversely, the backward method achieves substantial gains by exploring trajectories toward a predefined initial state, thus mitigating the overestimation risk associated with unknown states (yang2025rtdiff). Finally, demonstrating consistently superior and robust performance, our bi-directional framework, BiTrajDiff, empirically validates that by simultaneously synthesizing forward-future and backward-history trajectories, BiTrajDiff constructs novel, high-quality paths absent from the original dataset.

### v4

@@ 1 Introduction
Offline Reinforcement Learning (Offline RL) (Fujimoto et al., 2019) seeks to derive effective decision policies exclusively from pre-collected datasets, obviating the necessity for online exploration due to the inherent risks and costs associated with real-time interaction (Levine et al., 2020). This paradigm has demonstrated notable efficacy across various real-world applications, including robotic control (Qing et al., 2024; 2022, Liu et al., 2023) and power grid management (Chen et al., 2024, Xu et al., 2024). A primary challenge in offline RL is addressing the out-of-distribution (OOD) problem, stemming from the inherent divergence between the static offline dataset distribution and the learned policy behavioral distribution during training. Consequently, the direct application of conventional online RL methods inevitably leads to extrapolation error (Levine et al., 2020), characterized by the overestimation of value functions for OOD state-action pairs. To tackle the OOD problem, offline RL research emphasizes the development of techniques that constrain learned policies to remain within the distribution of the static offline dataset. These approaches include imposing explicit constraint terms during policy optimization (Qing et al., 2024, Chen et al., 2022) and penalizing the value function estimates for OOD actions (Lyu et al., 2022, Xu et al., 2023).
@@
Despite promising advancements in offline RL, current algorithms critically restrict agent actions, as evidenced by (Levine et al., 2020, Fujimoto et al., 2019, Qing et al., 2024). This limitation stems from the inherent mechanism constraining learned policy to the biased distribution of offline datasets, often leading to overfitting on limited behavioral patterns. While this principle effectively facilitates the replication of favorable behavioral patterns within the data, it inevitably curtails generalizability, hindering the exploration of optimal behaviors beyond dataset coverage. A common approach to mitigate the distribution bias is Data Augmentation (DA) method, which enriches the offline dataset with a large amount of synthesized high-fidelity data. For instance, studies like (Lu et al., 2023, Li et al., 2024, Li and Zhang, 2024) leverage one generative diffusion model to approximate the distribution of single-direction forward-future trajectories for augmenting offline data. Notably, as data augmentation (DA) methods operate by modifying the training dataset, they are orthogonal to specific offline RL algorithms, enabling them to enhance any such algorithm without altering its core optimization mechanisms. However, a critical limitation of DA methods is their exclusive focus on trajectories originating from a predefined initial state. While these generated single-direction trajectories effectively explore behaviors emerging from a given state, they neglect the exploration of trajectories leading to that state. This oversight restricts the utility of the generated trajectory to locally explore the trajectory space immediately surrounding existing data, specifically for transitions originating from states already in the dataset. Consequently, while DA methods successfully generate trajectories for local state pairs already reachable within the offline dataset, they are ineffective in discovering global trajectories for previously unreachable state pairs. This naturally leads us to a neglected yet unresolved question:
@@ 2 Related Works
Offline Reinforcement Learning encompasses four primary categories: policy constraint (Wang et al., 2022, Xu et al., 2023), value regularization (Kostrikov et al., 2021, Kumar et al., 2020), model-based (Yu et al., 2020, Chemingui et al., 2024), and return-conditioned supervised learning (Chen et al., 2021, Paster et al., 2022). Policy constraint methods restrict policies to the offline dataset, employing techniques such as explicit behavior cloning regularization (Qing et al., 2024, Chen et al., 2022) or implicit optimization towards in-sample optimal policies (Nair et al., 2020, Yue et al., 2022). Meanwhile, value regularization promotes conservative value functions to alleviate OOD overestimation via establishing precise lower bounds for Q-values to facilitate in-sample action selection (Lyu et al., 2022, Park et al., 2024). Model-based approaches construct task-oriented world models from offline data, including one-step dynamics (Yu et al., 2020, Chemingui et al., 2024) and multi-step sequence models (Hafner et al., 2020; 2023), which are subsequently used for offline reward penalization (Yu et al., 2021, Kidambi et al., 2020) and online decision planning to ensure in-distribution actions (Hansen et al., 2022; 2023). Return-conditioned supervised learning trains trajectory generators conditioned on returns, with transformers (Chen et al., 2021, Schmied et al., 2024) and diffusion models (Janner et al., 2022, Ajay et al., 2022) as common backbones. Despite their effectiveness, these approaches are fundamentally limited by reliance on the dataset distribution.
Data Augmentation for Offline RL synthesizes additional interactions to enrich the dataset for subsequent RL training. Early work employs dynamics ensembles to generate transitions with uncertainty weighting (Zhang et al., 2023) or diffusion models for flexible one-step oversampling (Lu et al., 2023). More recent efforts shift toward trajectory-level generation. Some work focuses on on-policy rollouts, guided either by policy outputs (Jackson et al., 2024) or interaction histories (He et al., 2023). Other approaches explore prioritized augmentation based on dynamics error or curiosity (Wang et al., 2024), as well as trajectory recomposition through reverse generation (Yang and Wang, 2025) or diffusion-based stitching (Li et al., 2024). Despite these advances, existing DA methods remain limited by their forward-only generation paradigm, restricting exploration to local variations around the dataset. Our BiTrajDiff overcomes this limitation by jointly generating forward and backward trajectories from intermediate states.
@@ 3 Preliminary
Reinforcement Learning paradigm (Sutton and Barto, 2018) is formally defined as a Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma,\rho_{0}\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the environment dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\rho_{0}$ denotes the initial state distribution. At each timestep $t$, an agent executes an action $a_{t}$ according to its policy $\pi(\cdot|s_{t})$ in the current state $s_{t}$, transitioning to the subsequent state $s_{t+1}$ with reward $r_{t}$ based on $P$. The objective of the agent is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return: $\pi^{*}=\arg\max_{\pi}\mathcal{J}(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. In contrast to online RL, offline RL (Levine et al., 2020) relys solely on a static trajectories dataset $\mathcal{D}=\{\tau^{i}\}_{i=1}^{N}$, where $N$ is the number of trajectories and each trajectory $\tau^{i}=\{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\}_{t=0}^{T-1}$ comprises a sequence of state-action-reward tuples.
Diffusion Models (Ho et al., 2020, Song et al., 2021) are generative frameworks that learn data distributions through a forward diffusion process and a reverse denoising process. The forward process gradually corrupts data $\mathbf{x}_{0}$ via a predefined noise schedule $\{\beta_{k}\}_{k=1}^{K}$, where each transition is defined as $q(\mathbf{x}_{k}|\mathbf{x}_{k-1})=\mathcal{N}\left(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}\right)$. As $K\to\infty$, the terminal distribution $q(\mathbf{x}_{K})$ converges to an isotropic Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process, parameterized by $\theta$, aims to iteratively reconstruct the original data through learnable Gaussian transitions: $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})=\mathcal{N}\left(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\Sigma_{\theta}(\mathbf{x}_{k},k)\right).$ By maximizing the Empirical Lower Bound (ELBO) (Sohn et al., 2015) on the log-likelihood of the sampled data, the diffusion model can be trained with a simplified surrogate loss (Sohn et al., 2015):
@@ 4.1 Bidirectional Diffusion Training
Building upon recent advances in diffusion-based sequence generation (Fathi et al., 2025, Gong et al., 2022), we propose a bidirectional trajectory generation framework comprising independent forward and backward trajectory diffusion models. These models are separately trained on offline datasets to learn the trajectory distribution of the behavior policy, facilitating bidirectional state trajectory generation from shared intermediate “anchor” states.
@@
where $\mathbf{x}_{0}(\tau)$ denotes the final generated subtrajectory of diffusion, $\mathbf{y}(\tau)$ represents the condition for generation, and $\theta$ represents the learnable parameters of the diffusion model. While various condition metrics are available, such as temporal difference error, Q-function (Wang et al., 2024), our framework utilizes the cumulative return $R$ and the state $s$ as conditional input $\mathbf{y}(\tau)$. This approach, consistent with prior works (Ajay et al., 2022, Ding et al., 2024), enables the generation of trajectories with desired returns starting from the specific states. However, to enable bidirectional trajectory generation, the formations of the backward and forward diffusion condition inputs are quite distinct.
@@
To realize conditional diffusion model training and inference, we employ the Classifier-Free Guidance (CFG) technique (Ho and Salimans, 2022), which uses a single denoise model to handle both conditional and unconditional generation. Formally, the perturbed noise after CFG can be formed as:
@@
where $\text{Bern}(p)$ represents the Bernoulli distribution. In our implementation, we fix the state condition $s_{t}$ in $\mathbf{x}_{k}(\tau)$ during the adding-noise training as well as the denoising inference stage for simplicity and efficiency following (Ajay et al., 2022, Yang and Wang, 2025), while the diffusion denoise models $\epsilon_{\theta}$ only take the cumulative return $R(\tau)$ components of $\mathbf{y}(\tau)$ as condition inputs.
@@ 4.2.1 Bidirectional State Trajectory Generation
We utilize the pretrained forward and backward trajectory diffusion model in section 4.1 to generate bidirectional state trajectories. Prior methods (Lu et al., 2023, Li and Zhang, 2024) model only the conditional distribution of trajectories originating from a given state, neglecting the historical pathways leading to those states. As a result, these data augmentation techniques yield only limited expansion in behavior diversity. To address this limitation without introducing additional noise, our proposed method, BiTrajDiff, generates both forward and backward trajectories and integrates them by stitching at a shared intermediate state $s_{t}$, which serves as a consistent anchor.
@@ 4.2.3 Bidirectional Trajectory Filter
In summary, our BiTrajDiff operates in two key stages. Initially, it trains a bidirectional diffusion model to capture the distributions of both forward-future and backward-history trajectories. Subsequently, BiTrajDiff iteratively generates bidirectional trajectory conditioned on intermediate anchor states sampled from $\mathcal{D}$. These generated sequences undergo further completion and filtering to create a high-fidelity augmented dataset $\tilde{\mathcal{D}}$, which is then mixed with $\mathcal{D}$ to train downstream offline RL algorithms. We adopt CleanDiffuser (Dong et al., 2024) as the backbone for BiTrajDiff, owing to its generality and robustness. The pseudocode of BiTrajDiff is presented in Appendix A.
@@ 5 Experiments
To demonstrate the effectiveness of the proposed BiTrajDiff method, we conduct experiments on the D4RL benchmark (Fu et al., 2020). Our evaluation seeks to answer the following questions: (1) Does BiTrajDiff outperform existing DA methods across various offline RL algorithms? (Section 5.2 and Appendix C) (2) How do different components of BiTrajDiff affect the offline RL performance? (Section 5.3 and Appendix D-F) (3) Can BiTrajDiff find the potentially valuable state leading to high return more effectively than other DA baselines? (Section 5.4 and Appendix G) (4) Can BiTrajDiff generate more diverse trajectories compared to single-direction DA approaches? (Section 5.5)
@@ 5.1 Experiment Settings
We evaluate our approach on the D4RL benchmark (Fu et al., 2020), focusing on locomotion, navigation, and manipulation tasks. We compare BiTrajDiff with three state-of-the-art DA methods: Synther (Lu et al., 2023),DiffStitch (Li et al., 2024) and RTDiff (Yang and Wang, 2025). To evaluate the effectiveness of them, we conduct extensive experiments on four representative offline RL algorithms: IQL (Kostrikov et al., 2021), TD3BC (Fujimoto and Gu, 2021), CQL (Kumar et al., 2020), and DT (Chen et al., 2021). These methods have been widely adopted as standard baselines in offline RL due to their stable performance. The hyperparameter setting can be viewed in Appendix B.2.
@@
Source Task  IQL (Kostrikov et al., 2021)  TD3BC (Fujimoto and Gu, 2021)
@@
Task  IQL (Kostrikov et al., 2021)  TD3BC (Fujimoto and Gu, 2021)
@@ 5.3 Ablation Analysis
Ablation Study of the Direction of Diffusion Generation. To confirm the effectiveness of our bidirectional generation paradigm, we compare the performance improvement with the augmented diffusion trajectories generated in a single direction and in a bidirectional way, as shown in Figure 2. While the forward method yields moderate improvements, it is consistently outperformed because it neglects the historical transitions leading to critical, high-reward states. Conversely, the backward method achieves substantial gains by exploring trajectories toward a predefined initial state, thus mitigating the overestimation risk associated with unknown states (Yang and Wang, 2025). Finally, demonstrating consistently superior and robust performance, our bi-directional framework, BiTrajDiff, empirically validates that by simultaneously synthesizing forward-future and backward-history trajectories, BiTrajDiff constructs novel, high-quality paths absent from the original dataset.